# DriveGPT4: Interpretable End-to-end Autonomous Driving via Large Language Model

### v4

@@ 5.2 End-to-end Control
| ✓ | ✓ | ✓ | 102.71 | 19.00 | 45.10 | 56.34 | 22.94 | 81.62 | 1.30 | 30.09 | 79.92 | 8.98 | 59.23 | 79.59 |
5.3 Qualitative Results.
Multiple qualitative results are provided for intuitive comparison. For concise visualization, we only show four frames of the input video clip. First, an example from the BDD-X testing set is visualized in Fig. 3. DriveGPT4 can generate high-quality texts and control predictions based on the prompt. Then, to verify the generalization ability of DriveGPT4, we apply DriveGPT4 to the NuScenes dataset (Caesar et al., 2020) for zero-shot QA in Fig. 4. We also try DriveGPT4 on video games to further test its generalization ability. An example is illustrated in Fig. 5.
@@ 5.4 Ablation Studies
In this paper, several ablation studies are conducted to validate proposed designs, and the results are provided in Tab. 8. By removing either BDD-X QAs or ChatGPT QAs during finetuning, a decrease in corresponding performance is observed, highlighting the significance of including all task-specific multimodal data. QA pairs generated by ChatGPT enable DriveGPT4 to answer human questions in more flexible patterns, and enhance the QA ability of BDD-X questions. Then, we test DriveGPT4 without the mix-finetune strategy by removing the general image and video instruction-following data. Severe performance deduction is observed, indicating the necessity of finetuning DriveGPT4 with diverse multimodal data. Thus, changes to DriveGPT4 would negatively impact its versatile QA capabilities for interpretable end-to-end autonomous driving.
6 Conclusion
@@
This paper presents DriveGPT4, an interpretable end-to-end autonomous driving system using multimodal LLM. A new dataset for autonomous driving interpretation is developed with the assistance of ChatGPT and employed to mix-finetune DriveGPT4, enabling it to respond to human inquiries about the vehicle. DriveGPT4 utilizes input videos and texts to generate textual responses to questions and predict control signals for vehicle operation. It outperforms baseline models in various tasks such as vehicle action description, action justification, general question answering, and control signal prediction. Moreover, DriveGPT4 exhibits generalization ability through zero-shot adaptation. In the future, DriveGPT4 will be further enhanced and adapted to close-loop vehicle control tasks.
Appendix A Data Processing
In the data processing part, we generate three question sets for fixed question answering, i.e., for action description, for action justification and for control signals. The detailed question sets are shown in Tab. 9, Tab. 10 and Tab. 11, respectively.
The detailed prompt to generate conversations by ChatGPT is shown in Tab. 12.
Appendix B Evaluation Scores Generated by ChatGPT
ChatGPT-generate QAs have long and complicated texts, thus ChatGPT score is needed. The prompt for evaluation score generation is demonstrated in Tab. 13. For each ChatGPT-generated QA pair, we embed them in the prompt texts and send them to ChatGPT. ChatGPT first outputs a numerical number ranging from 0 to 1, and then provides explanations to the score. Since ChatGPT score is not stable, we report the mean of 3 rounds of evaluations for qualitative comparison.
Differently, BDD-X QAs have a short and relatively fixed format. We did not use ChatGPT score to evaluate BDD-X QAs since ChatGPT score is quite unstable and unpredictable. For BDD-X QAs evaluation, compared with ChatGPT score, conventional metrics like CIDEr (Vedantam et al., 2015) utilized in the NLP community are more reliable.
Appendix C Prediction Consistency
| Method | ChatGPT |
| ADAPT | 79.40 |
| DriveGPT4 | 80.69 |
To measure whether DriveGPT4 makes consistent control signal and text response predictions, ChatGPT is utilized for consistency evaluation. ChatGPT is prompted to generate a score ranging from 0 to 1 to measure the consistency between generated text answers and predicted control signals. The detailed prompt for ChatGPT is demonstrated in Tab. 14. Consistency evaluation examples are shown in Fig. 7. The evaluation results are listed in Tab. 15. The consistency scores of ADAPT and DriveGPT4 are similar. In most situations, both ADAPT and DriveGPT4 can have consistent predictions.
Appendix D Comparison with GPT4-V
As the multimodal version of GPT4, GPT4-V can understand and reason single-frame images and illustrate excellent generalization ability for various daily tasks. However, GPT4-V is still a general model for images, and not specially finetuned for grounding autonomous driving applications. Before the comparison, GPT4-V is prompted with several BDD-X QA pairs in advance. During the qualitative evaluation, even though GPT4-V illustrates powerful recognition and reasoning ability, it is observed that it (1) cannot predict numerical control signals; (2) fails to correctly understand some vehicle actions, especially dynamic actions (e.g., turning, accelerating, etc.). Examples of the comparison between GPT4-V and DriveGPT4 are shown in Fig. 8, Fig. 9 and Fig. 10.
Appendix E Additional Qualitative Results
We provide more qualitative results of DriveGPT4 on the BDD-X dataset for interpretable end-to-end autonomous driving in Fig. 11 to 14.

### v5

@@ IV-B Training
| ✓ | ✓ | ✓ | 99.10 | 18.32 | 44.73 | 56.34 | 22.94 | 81.62 | 1.30 | 30.09 | 79.92 | 8.98 | 59.23 | 79.59 |
V Experiment
V-A Interpretable Autonomous Driving
In this section, we evaluate DriveGPT4 and its baselines on interpretation generation, covering vehicle action description, action justification, and additional questions about vehicle status. ADAPT [14] serves as the state-of-the-art baseline work. Recent multimodal video understanding LLMs [40, 39] are also considered for comparison. All methods use 8-frame videos as input. Currently, DriveGPT4 does not take 32-frame videos as input like ADAPT considering the heavy memory consumption and inference speed, which could be treated as a limitation of this work.
Testing Set Split. During vehicle driving, the distribution of scenes is usually not balanced. For example, some simple scenes like driving straight-forward are more commonly seen than more challenging vehicle turning or lane changes. For a comprehensive evaluation comparison, the testing set is split into “Easy”, “Medium” and “Hard” sets based on the driving scene and vehicle status. Detailed split information is shown in Tab. III.
Evaluation Metrics. To thoroughly assess the methods, we report multiple metric scores widely used in the NLP community, including CIDEr [50], BLEU4 [51], and ROUGE-L [52]. The BDD-X QA task tends to have a fixed format, so the aforementioned NLP metrics are already sufficient for evaluation. ChatGPT-generated QAs possess flexible formats and more complicated semantic meanings. Following past MLLM works [27, 42, 39], we also report the score generated by ChatGPT. ChatGPT is prompted to assign a numerical score between 0 and 1, with a higher score indicating better prediction accuracy. The detailed prompt for ChatGPT-based evaluation is available in the appendix. However, it should be noted that the ChatGPT score is not stable, thus we report the mean of three times of evaluations for reference.
Action Description and Justification. The goal is to predict vehicle action descriptions and justifications as closely as possible to the given labels. Evaluation results of all testing splits are displayed in Tab. IV. More detailed results are shown in Tab. V. From the results, it is observed that DriveGPT4 outperforms the previous SOTA baseline ADAPT on all testing data, especially for the “Hard” splits with more challenging driving scenes and vehicle dynamics. The effectiveness and superiority of the proposed DriveGPT4 are well demonstrated.
Additional Question Answering. The above vehicle action description and justification have relatively fixed formats. To further evaluate the interpretable ability and flexibility of DriveGPT, additional questions are generated following section III-A. A hundred randomly sampled video clips in the BDD-X testing set are used for question generation. Compared with action descriptions and justifications, these questions are more diverse and flexible. The evaluation results are shown in Tab. VII. ADAPT cannot answer additional questions except for the vehicle action description and justification. Previous video understanding LLMs can answer these questions but they do not learn autonomous driving domain knowledge. Compared with all baselines, DriveGPT4 presents superior results, demonstrating its flexibility.
V-B End-to-end Control
In this section, we evaluate DriveGPT4 and its baselines for open-loop control signal prediction, specifically focusing on speed and turning angle. All methods are required to predict control signals for the next time step. Following previous works on control signal prediction, we use root mean squared error (RMSE) and threshold accuracies () for evaluation. measures the proportion of test samples with prediction errors lower than . For a comprehensive comparison, we set with multiple values: . The quantitative results for the previous state-of-the-art (SOTA) method ADAPT and DriveGPT4 are shown in Tab. VI. DriveGPT4 achieves superior results for both speed and turning angle predictions.
V-C Qualitative Results.
Multiple qualitative results are provided for intuitive comparison. For concise visualization, we only show four frames of the input video clip. First, an example from the BDD-X testing set is visualized in Fig. 3. DriveGPT4 can generate high-quality texts and control predictions based on the prompt. Then, to verify the generalization ability of DriveGPT4, we apply DriveGPT4 to the NuScenes dataset [53] for zero-shot QA in Fig. 4. We also try DriveGPT4 on video games to further test its generalization ability. An example is shown in Fig. 5.
@@ V-D Ablation Studies
In this paper, several ablation studies are conducted to validate proposed designs, and the results are provided in Tab. VIII. By removing either BDD-X QAs or ChatGPT QAs during finetuning, a decrease in corresponding performance is observed, highlighting the significance of including all task-specific multimodal data. QA pairs generated by ChatGPT enable DriveGPT4 to answer human questions in more flexible patterns, and enhance the QA ability of BDD-X questions. Then, we test DriveGPT4 without the mix-finetune strategy by removing the general image and video instruction-following data. Severe performance deduction is observed, indicating the necessity of finetuning DriveGPT4 with diverse multimodal data. Thus, changes to DriveGPT4 would negatively impact its versatile QA capabilities for interpretable end-to-end autonomous driving.
VI Conclusion
@@
This paper presents DriveGPT4, an interpretable end-to-end autonomous driving system using multimodal LLM. A new dataset for autonomous driving interpretation is developed with the assistance of ChatGPT and employed to mix-finetune DriveGPT4, enabling it to respond to human inquiries about the vehicle. DriveGPT4 utilizes input videos and texts to generate textual responses to questions and predict control signals for vehicle operation. It outperforms baseline models in various tasks such as vehicle action description, action justification, general question answering, and control signal prediction. Moreover, DriveGPT4 exhibits generalization ability through zero-shot adaptation. In the future, DriveGPT4 will be further enhanced for close-loop vehicle control tasks. To handle the drifting issue of imitation learning, an LLM expert will be developed for data collection without human effort.